# Network Formation and Dynamics Among Multi-LLMs

### v3

@@ Acknowledgements.
Appendix A Regression Table
In Table S.1, we report the regression coefficient for the regression in the real-world network data. The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
@@ Appendix A Regression Table
Table S.1. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown in parentheses. The null hypothesis corresponds to the respective parameter being equal to 0. We report the percent change in accuracy, average path length, and average clustering coefficient compared to the initial network (before the deletion of edges). For the change in modularity, we run the Louvain algorithm ten times and perform a t-test with the resulting modularities. For the UChicago30 dataset, we report the t-statistic value in the subgraph induced by the 2,000 sampled nodes, since the newly added edges would have a very small effect on the change in the community structure if we were to measure it in the whole network. We also report the modularity change (t-statistic) of the whole graph inside brackets.
Appendix B Analytical Regression Tables
@@
Table S.2. Multinomial logit coefficients for Caltech36. The standard errors of the estimates are shown in parentheses.
@@
Table S.3. Multinomial logit coefficients for Swarthmore42. The standard errors of the estimates are shown in parentheses.
@@
Table S.4. Multinomial logit coefficients for UChicago30. The standard errors of the estimates are shown in parentheses.
Appendix C LLM Reasoning Summarization
In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions. After LLMs form networks, we ask them about their reasoning. We then use GPT models (GPT-3.5 for Section 1.A. and GPT-4 for Section 1.C) to group their explanations into common themes, creating various categories. To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
C.1. Micro-level Principles
In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 1.A.
We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases. This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment.
For Figure 7(b), we observe an opposite pattern in the LLM reasoning.
A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
Finally, Figure 7(c) contains the reasons for Figure 3 based on having a common hobby, a common color, or a common location. We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
(a) Reasons for Figure 1(a)
(b) Reasons for Figure 2(a)
(c) Reasons for Figure 3(b)
Figure S.7. Distribution of reasoning across categories for the networks of Figures 1(a), 2(a), and 3. The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies). The error bars correspond to 95% confidence intervals.
C.2. Real-World Data
We also present the LLM reasoning for real-world networks. The reasons have been set so that there is a one-to-one correspondence with the regression coefficients. Specifically, the Mutual Friends reason corresponds to $\hat{\theta}_{\mathrm{TC}}$, the Number of friends reason to $\hat{\theta}_{\mathrm{PA}}$, and the Similar attributes reason corresponds to $\hat{\theta}_{\mathrm{H}}$ respectively. Consistently with the results reported in Table S.1, the agents’ reasoning is more geared towards homophily (same attributes), then triadic closure (mutual friends), and finally preferential attachment (number of friends). These results are consistent with the regression.
Figure S.8. Distribution of reasoning across categories for the real-world networks. The classification is done using GPT-4 on a set of 20 randomly sampled reasons repeated over 5 times. In agreement with Table S.1, the multi-LLMs put more weight on homophily, then on triadic closure, and then on preferential attachment. The error bars correspond to 95% confidence intervals.
Appendix D Network Evolution
Here we depict the evolution of the networks generated by the LLM agents.
D.1. Principle 1: Preferential Attachment
We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes. We observe that for the temperature being 0.5 we have a core-periphery like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
@@ D.1. Principle 1: Preferential Attachment
D.2. Principle 2: Triadic Closure
@@
Appendix E Feature Representations for Prompts
E.1. Principle 1: Preferential Attachment
For the structure-based information (Figure 1(a)) we have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 1(b)) we have the following features:
"name" : 0,
"degree" : 4
E.2. Principle 2: Triadic Closure
For the structure-based information (Figure 2(a)) we have the following features:
"name" : 0,
"common_neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 2(b)) we have the following features:
"name" : 0,
"num_common_neighbors" : 4
E.3. Principle 3: Homophily
We have the following features:
"name" : 0,
"favorite_color" : "red",
"hobby" : "hiking",
"location" : "Boston"
E.4. Principle 5: Small-World
We have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
E.5. Real-World Data
We have the following features:
"name" : 0,
"status" : "student",
"major" : 10,
"second major" : 93,
"accommodation" : "house",
"high_school" : 5,
"graduation_year" : 2008
We note that the initial Facebook100 dataset included gender information as a feature. We chose not to include the gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019).

### v4

@@ Acknowledgements.
Appendix A Full Regression Table for GPT-4 (gpt-4-1106-preview) and the Facebook100 Data
In Table S.2, we report the regression coefficient for the regression in the real-world network data for all temperatures and GPT-4 (gpt-4-1106-preview). The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
@@ Appendix A Full Regression Table for GPT-4 (gpt-4-1106-preview) and the Facebook100 Data
Table S.2. Multinomial logit coefficients for three networks from the Facebook100 dataset and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses. The null hypothesis corresponds to the respective parameter being equal to 0. We report the percent change in accuracy, average path length, and average clustering coefficient compared to the initial network (before the deletion of edges). For the change in modularity, we run the Louvain algorithm ten times and perform a t-test with the resulting modularities. For the UChicago30 dataset, we report the t-statistic value in the subgraph induced by the 2,000 sampled nodes, since the newly added edges would have a very small effect on the change in the community structure if we were to measure it in the whole network. We also report the modularity change (t-statistic) of the whole graph inside brackets.
A.1. Analytical Regression Tables for GPT-4 (gpt-4-1106-preview)
@@
Table S.3. Multinomial logit coefficients for Caltech36 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
@@
Table S.4. Multinomial logit coefficients for Swarthmore42 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
@@
Table S.5. Multinomial logit coefficients for UChicago30 and GPT-4 (gpt-4-1106-preview). The standard errors of the estimates are shown in parentheses.
Appendix B Network Evolution and Omitted Simulations
Here we depict the evolution of the networks generated by the LLM agents, as well as omitted simulations.
B.1. Principle 1: Preferential Attachment
Network Evolution
We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes. We observe that for the temperature being 0.5 we have a core-periphery-like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
@@ B.1. Principle 1: Preferential Attachment
Simulations with Degree Information
In Figure S.9 we provide the results with degree-information only. We observe that the agents form connections around high-degree nodes only (see Figure S.9). The same result (star-like networks) holds for the other LLM models and temperatures.
Figure S.9. Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures. For the degree-based simulations, node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$). With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
B.2. Principle 2: Triadic Closure
Network Evolution
@@ B.2. Principle 2: Triadic Closure
Simulations with the Number of Common Neighbors
Instead of giving the neighborhood information, the simulations presented in Figure S.11 use the number of common neighbors. We observe behavior similar to Figure 2.
Figure S.11. Results for Principle 2 (triadic closure). The figure shows the same networks as in Figure 2 with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). The error bars correspond to 95% confidence intervals.
Appendix C Chain-of-Thought Experiments
# Task You are located in a school. Your task is to select a set of people to be friends with. # Profile Your profile is given below after chevrons: <PROFILE> { "name" : "Person 0", "favorite subject" : "Chemistry", "neighbors" : ["Person 3", "Person 432", "Person 4", "Person 3", "Person 32"] } </PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES> [ { "name" : "Person 1", "favorite subject" : "Mathematics", "neighbors" : ["Person 3", "Person 4", "Person 23", "Person 65"] }, { "name" : "Person 33", "favorite subject" : "History", "neighbors" : ["Person 342", "Person 2", "Person 12"] }, ... ] </PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "reason" : reason for selecting the person, "name" : name of the person you selected }}, ... ] # Notes - The output must be a list of JSON objects ranked in the order of preference. - You can make at most 1 selection.
ALGORITHM 3 Example prompt regarding social network data with Chain-of-Thought reasoning. Note that compared to Algorithm 1 the order of the fields name and reason in the output format is reversed.
We experiment with Chain-of-Thought (CoT) reasoning (Wei et al., 2022). To induce CoT reasoning we ask the LLM agents to output the reason and then their choice (i.e. by reversing the order of reason and name in Algorithm 2). The resulting prompt can be found at Algorithm 3. In the following figures, we show the results from the same experiments as the ones we of the main text with the different that CoT is used.
(a) Probability of connecting to top-$k$ nodes for different models, temperatures, and environments
(b) Power law fits ($\hat{\gamma}$) and standard errors for different models, temperatures, and environments
Figure S.12. Results for Principle 1 with CoT reasoning (preferential attachment) The multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$. Top: Probability of connecting to top-$k$-degree nodes for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-3.5 and environment to baseline) and environment (model fixed to GPT-3.5 and environment temperature to 1.5) for networks generated according to Principle 1 with $n=200$ nodes. Bottom: Power Law exponents and standard errors for varying model, temperature, and environment.
(a) Probability of connecting to top-$k$ for different models, temperatures, and environments
(b) Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for different models, temperatures, and environments
Figure S.13. Results for Principle 2 with CoT reasoning (triadic closure). Top: Probability of connecting to top-$k$ nodes (in terms of common neighbors) for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-4 Mini and environment to baseline) and environment (model fixed to GPT-4 Mini and environment temperature to 0.5) for networks generated according to Principle 2 ($n=50$, 10 simulations for each model, environment and temperature). Bottom: Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for networks generated according to Principle 2 in different models, temperatures, and environments.
(a) Assortativity and Louvain Modularity with different LLM models and environments
Figure S.14. Results for Principle 3 (Homophily) and Principle 4 (Community structure due to homophily) with CoT reasoning. Top: Assortativities and Louvain modularity according to Principle 3 ($n=50$, 5 simulations for each row) in different environments (school, work, community) using different models. The statistical significance is $P<0.001$ for all t-tests (comparing with 0).
(a) Regression plot for different models and environments for $\beta=0.25$ and $k=5$.
Figure S.15. Fitted results for Principle 5 with CoT reasoning (small world). Regression plot for the relation $L\sim\log(n)$ for different LLM models for $\beta=0.25$ and $k=5$. The legend shows the effect size ($a$) and the $P$-value. (*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.)